OpenStudy (anonymous):

find the equation of the tangent aat x=3 to the curve with the equation y=2x^2-3x+2.

OpenStudy (phi):

did you find the derivative of y= 2x^2-3x +2 dy/dx will be the slope of the curve, after you sub in x=3 into it also, when x=3, y= 2*3^2- 3*3 +2 = 11 so you will have a point (3,11) and the slope use y= mx+b sub in the slope for m, and replace x,y with 3,11 and solve for b

OpenStudy (anonymous):

is the derivative 4x-3+0?

OpenStudy (anonymous):

That is the correct derivative.

OpenStudy (phi):

yes dy/dx = 4x-3 (people don't bother to write +0) now find m m= dy/dx with x=3

OpenStudy (anonymous):

how do you find m?

OpenStudy (anonymous):

@emzy_777 just put the value of x in dy/dx:)

OpenStudy (anonymous):

i.e. at x=3 dy/dx=4(3)-3=12-3=8 so y=mx+c y=8x+c it must pass through (3,11) so 11=8(3)+c c=11-24=-13 so final equation is y=8x-13:) @emzy_777

OpenStudy (anonymous):

thank you @Aperogalics:)

OpenStudy (anonymous):

ur welcome:)

OpenStudy (anonymous):

@Aperogalics i just checked the answer and the final equation was 9x-7=16...:/

OpenStudy (anonymous):

sorry *9x-y=16

OpenStudy (anonymous):

ya my fault at x=3 dy/dx=4(3)-3=12-3=9 not 8:( so y=9x+c 11=9(3)+c c=11-27 c=-16 :)

OpenStudy (anonymous):

aha its cool:) thank youuu